If the perimeter of a rectangle is 34 ft and the width is 8 ft, find its length. [Hint: The perimeter of a rectangle is given by P = 2L + 2W.]
Answer by jojo14344(1513)

Direct substitution, by P=2l+2W
34=2L+2(8)
34-16=2L
L=18/2= 9ft